aspect0943:

Part A: Eveline rented a car at $180 for 4 days. If she rents the same car for 9 days, she has to pay a total rent of $325. Write an equation in the standard form to represent the total rent (y) that Eveline has to pay for renting the car for x days. (4 points) Part B: Write the equation obtained in Part A using function notation. (2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

aspect0943:

i got 30 minutes

TutorJohnfoster:

Part A (4,180) (9,325) slope =\[\frac{ y2-y1 }{ x2-x1 }\] slope = \[\frac{ 325-180}{9-4 }\] slope =29 y=mx+c where m is the slope Using one of the point 325=29(9)+c 325=261+c c=325-261 c=64 y=29x+64 Part B f(x)=29x+64 Part C To graph the above equation you just need to plot the points (4,180) and (9,325) and then draw a line joining the two points.